Blake Inc., a U.S. MNC, needs to pay EUR1,000,000 in one year. It can earn 2 percent annualized on a German security. The current spot rate for the euro is USD1.00 per euro. Blake can borrow funds in the U.S. at an annualized interest rate of 0 percent. If Blake uses a money market hedge to hedge the payable, what is the cost of implementing the hedge (rounded to the nearest dollar)?

$980,392

None of the answers is correct.

$1,000,000.

$19,608.
Answer #1

Amount payable, FV = Euro 1,000,000

Interest rate in Euro = r = 2%

Hence, present value of the payable = PV = FV / (1 + r)t = Euro 1,000,000 / (1 + 2%)1 = Euro  980,392

Spot rate, S =  USD 1.00 per euro

Hence, present value of payable in USD = PV x S =  Euro 980,392 x USD 1.00 per Euro = USD 980,392

Hence, the correct answer is the first option showing $ 980,392